### v1

# M2R2: MulitModal Robotic Representation for Temporal Action Segmentation

Abstract
Temporal action segmentation (TAS) has long been a key area of research in both robotics and computer vision. In robotics, algorithms have primarily focused on leveraging proprioceptive information to determine skill boundaries, with recent approaches in surgical robotics incorporating vision. In contrast, computer vision typically relies on exteroceptive sensors, such as cameras. Existing multimodal TAS models in robotics integrate feature fusion within the model, making it difficult to reuse learned features across different models. Meanwhile, pretrained vision-only feature extractors commonly used in computer vision struggle in scenarios with limited object visibility. In this work, we address these challenges by proposing M2R2, a multimodal feature extractor tailored for TAS, which combines information from both proprioceptive and exteroceptive sensors. We introduce a novel pretraining strategy that enables the reuse of learned features across multiple TAS models. Our method achieves state-of-the-art performance on the REASSEMBLE dataset, a challenging multimodal robotic assembly dataset, outperforming existing robotic action segmentation models by 46.6%. Additionally, we conduct an extensive ablation study to evaluate the contribution of different modalities in robotic TAS tasks.
I Introduction
Temporal action segmentation (TAS) is a long-standing challenge in both robotics and computer vision. The goal is to identify temporal boundaries between actions in untrimmed data and assign action labels to each segment, depending on the subsequent tasks. In robotics, TAS helps break down complex demonstrations into simpler actions for learning motion policies [1], or action preconditions and effects [2]. In computer vision, TAS is applied to tasks like video surveillance or medical procedure documentation [3].
Figure 1: The overview of mutimodal temporal action segmentation (TAS) using the proposed M2R2 representation, when applying to contact-rich assembly and disassembly task.
In the robotics community, researchers typically leverage proprioceptive sensory information from the robot, such as forces, torques, joint positions, and end-effector poses and twists, to solve the TAS task. These data streams are analyzed to identify meaningful boundary points that indicate action transitions. Existing approaches often rely on heuristics, such as detecting change points in force and torque derivatives [4], identifying sections of linear motion [1], using handcrafted feature spaces combined with SVM classifiers [5], or recognizing predefined skills in the demonstrations [6]. In recent years, the field of surgical robotics, has focused on leveraging vision and kinematic data for TAS, a task often referred to as surgical gesture recognition [7].
On the other hand, the computer vision community primarily focuses on human temporal action segmentation, where proprioceptive data is unavailable. As a result, existing methods rely on exteroceptive sensors such as videos [8], optical flow [8], or object accelerations [9] as model inputs. Pretrained or fine-tuned feature extractors are then used to obtain meaningful latent representations of the observations, which serve as inputs to multi-stage prediction models. These models iteratively refine action and boundary predictions at each stage [8]. Prior research has shown that deep learning-based TAS models can effectively learn meaningful feature representations and decompose complex, long-horizon demonstrations into simpler action sequences [8, 10, 11].
Despite promising advances in leveraging vision and multimodal data for robotic TAS in surgical robotics, proprioceptive-based approaches remain the predominant choice for larger scale manipulation tasks such as assembly and disassembly. However, these methods often rely on handcrafted heuristics, requiring extensive hyperparameter tuning or the construction of a predefined skill library before segmentation [6, 4]. In contrast, purely vision-based approaches struggle in scenarios with occlusions or visually similar objects. For example, distinguishing between two pegs that differ by only a few millimeters is particularly challenging due to their small size, yet such precision is crucial for learning accurate insertion policies. While multimodal approaches have shown promise in surgical robotics, they are typically trained end-to-end, restricting flexibility in modifying the action segmentation head. This limitation makes it difficult to incorporate state-of-the-art TAS models or adapt to new models as they emerge. In contrast, separating multimodal feature fusion and extraction from the TAS model allows for seamless integration of improved TAS architectures in the future.
To address these channels, in this work, we develop a multimodal deep learning feature extractor and a pretraining strategy tailored for robotic temporal action segmentation tasks, which is agnostic of the TAS model itself, which foster easier resuability of the learned features. We refer to the resulting fused features as multimodal robotic representation (M2R2) features. To compute the M2R2 features, we leverage proprioceptive information in the form of force and torque measurements at the robot’s wrist, end-effector pose and velocity, and gripper width. For exteroceptive data, we use RGB cameras and audio. In this work, we adopt a late fusion strategy, where each modality is first processed independently and then fused using a transformer-based model [12]. In our experiments, we show that the resulting M2R2 features improve temporal action segmentation performance over existing approaches on the REASSEMBLE dataset [13] and outperform using proprioceptive or exteroceptive information alone. Moreover, we demonstrate that these pretrained features can be used with existing deep learning-based temporal action segmentation models to successfully perform robotic temporal action segmentation. A general overview of the proposed approach is shown in Figure 1.
To summarize, our main contributions are as follows:
  1. 1.
    A deep-learning-based multimodal feature extractor for robotic temporal action segmentation in contact-rich manipulation tasks.
  2. 2.
    A pretraining strategy for learning multimodal features for robotic temporal action segmentation.
  3. 3.
    An extensive evaluation of the influence of different sensor modalities on robotic temporal action segmentation performance.
II Related works
II-A Deep Learning TAS
Many deep learning TAS models use pretrained visual features instead of training end-to-end [8, 10, 11]. These feature extractors must capture both short- and long-term dependencies in the data. Over the years, several video feature extractors have been proposed. Carreira et al. introduced I3D features [14], leveraging both RGB and optical flow in a two-stream 3D convolutional network. More recently, Wang et al. introduced ActionCLIP [15], which uses a contrastive text-to-video pretraining strategy to extend CLIP image features to video classification. However, these features are not well-suited for TAS, as they were trained on trimmed action clips. To address this, Li et al. developed BR-Prompt [16], by fine-tuning ActionCLIP to understand action execution order. Other approaches, such as ASPNet [9], fuse multiple modalities (images and object acceleration) using separate subspaces for each modality, with a shared subspace for information exchange.
In the second stage, TAS models are trained on the extracted features. Farha et al. introduced MSTCN [8] to reduce over-segmentation by refining predictions across multiple stages. Alternatively, Ishikawa et al. proposed the ARSF model [10], which regresses temporal action boundaries in addition to per-frame class predictions. More recently, models have incorporated modern architectures, including diffusion models [11]. Additionally, some approaches refine temporal segmentation by performing hierarchical refinement on the predicted action segments [17].
While much of the deep learning TAS community has focused on human action segmentation, where proprioceptive data is typically unavailable, robotic TAS benefits from proprioceptive data, which can significantly improve action segmentation performance. In this work, we demonstrate how techniques used to learn robust video feature extractors can be extended to fuse both proprioceptive and exteroceptive data, leading to better segmentation results. Furthermore, we show that modern TAS models can effectively leverage these multimodal features, seamlessly integrating into existing robotic TAS frameworks without requiring architectural modifications.
II-B Robotics TAS
In recent years, surgical robotics has increasingly focused on vision-based and multimodal TAS, also known as surgical gesture recognition, he goal is to identify actions performed by a human-controlled surgical robot. Earlier methods used Convolutional Neural Networks (CNNs) for action segmentation from vision data [18]. More recent approaches employ transformer models or CNNs for proprioceptive data [19], or combine proprioceptive and kinematic data [20]. However, existing multimodal TAS models merge feature extraction, and temporal action segmentation into a single end-to-end model, limiting the reusability of learned feature extractors across different TAS architectures. In contrast, decoupling feature extraction from the TAS model allows the learned features to be reused with various TAS models.
Despite the success of multimodal TAS in surgical robotics, proprioception-only TAS remains the primary method for large-scale robotic manipulation tasks, such as assembly and disassembly. Proprioception-based TAS can be classified into supervised and unsupervised methods. Supervised methods rely on labeled data to model primitive actions for recognizing actions in new task demonstrations. Early approaches [6] assume access to skill library, while more recent methods segment skills by comparing the current scene state to a symbolic representation of known skills together with a Support Vector Machine (SVM) to classify skills within a sliding window using handcrafted features like power, and contact duration [5].
Figure 2: M2R2 Model Architecture. To compute the multimodal feature at time instant $t_{i}$, we first process each modality separately to obtain image features $I_{i}$, audio features $A_{i}$, and proprioceptive features $\{S_{i}^{s}\}_{s=1}^{N_{s}}$, which are later fused using a Transformer encoder layer followed by an MLP. To obtain $I_{i}$, we use the ActionCLIP image encoder [15]. For $A_{i}$, we extract features using the Audio Spectrogram Transformer [21]. For the proprioceptive data, we compute $S_{i}^{s}$ by first upscaling the raw sensory data through a linear projection into a higher-dimensional space using a learnable projection matrix $W^{s}_{p}$. Next, we embed temporal information by applying an element-wise multiplication with a learnable temporal embedding matrix $E^{s}_{t}$. Finally, we compute the average over the temporal dimension to obtain $S_{i}^{s}$.
Unsupervised methods detect action boundaries without predefined skill models. Probabilistic approaches use Gaussian Mixture Models (GMMs) [22] or HMMs [23] to infer segmentation points. More recent methods, such as [4], use Bayesian Online Changepoint Detection to identify transitions in noisy force and torque data. Other approaches model action segments as linear trajectories, detecting boundaries based on reconstruction error relative to a predefined threshold [1].
Despite these advances, current TAS models in robotics have limitations. Many rely on predefined skill libraries, requiring extensive prior knowledge and reducing generalizability [6, 5]. Others depend on handcrafted features for action identification, which may not scale to diverse tasks [20]. Furthermore, many approaches need extensive hyperparameter tuning, making them difficult to adapt to new environments [4, 1]. In contrast, M2R2 directly processes raw sensory data using a deep neural network feature extractor that learns the most relevant representations for TAS. By decoupling feature extraction from TAS, M2R2 improves reusability and allows seamless integration with various TAS models.
III Methodology
III-A Problem formulation
The goal of our work is to learn a multimodal feature extractor $F_{m}$ for temporal action segmentation that fuses both exteroceptive and proprioceptive data. Given a set of visual observations $V=\{v_{i}\}_{i=1}^{T_{image}}$, audio observation $A=\{a_{i}\}_{i=1}^{T_{audio}}$, and proprioceptive observation $P$, we aim to extract multimodal features $X=\{x_{i}\}_{i=1...T}$ such that $X=F_{m}(V,A,P)$, where
$P=\bigcup_{k=1}^{N_{s}}\left\{\left(t_{k,i},s_{k}(t_{k,i})\right)\mid t_{k,i}\in T_{k}\right\}$ (1)
represents the collection of $N_{s}$ proprioceptive sensor modalities, with $t_{k,i}$ denoting the timestamp of the $i$-th measurement from sensor $k$, and $T_{k}$ representing the set of measurement timestamps for sensor $k$. Since different sensors operate at varying frequencies, we fuse the extracted features such that the resulting sequence aligns with the slowest sensor, which, in our case, is the camera. Finally, we use the extracted features $X$ to predict action segmentation labels $C=\{c_{i}\}_{i=1}^{T}$ by applying a state-of-the-art deep learning TAS model, $F_{seg}$, such that $C=F_{seg}(X)$. Figures 2 and 3 illustrate the process of learning the feature extractor $F_{m}$. Figure 2 shows how information from different modalities is fused at a single time step $i$ to obtain $x_{i}$, while Figure 3 depicts how the extracted features are processed for pretraining the M2R2 feature extractor.
III-B Preprocessing
Generally, different sensors operate at varying frequencies. Standard RGB cameras work at a relatively low frequency of around 10–60 Hz, while most robot proprioceptive data can be read at rates between 100 Hz and 1 kHz. Microphones, on the other hand, record at frequencies above 16 kHz. Simply selecting the samples of the higher-frequency data at the slowest sensor timestamps is insufficient, as it may result in the loss of valuable high-frequency information. For example, in contact-rich manipulation tasks, impacts and contact transitions cause rapid fluctuations in force readings, which could be lost if the measurements are downsampled to a very low frequency. Similarly, resampling audio signals to lower rates could result in the loss of higher-pitched sounds, such as clicking noises. To address this, we employ the following preprocessing procedure. Given a time instant $t_{i}$ at which a camera frame is published and the previous frame’s timestamp $t_{i-1}$, we collect prioproceptive sensor measurements within the window $[t_{i-1},t_{i})$. Next, we resample the collected data to a common frequency $f$, which is lower than the sampling rate of all sensors, by using linear interpolation. In our case, we select $f=300$ Hz. Additionally, for each sensor, we compute the mean and standard deviation over the entire training dataset, which we later use to normalize the readings.
For audio signals, we first resample the sensor readings within the window to 16 kHz. Following previous works, we compute the audio spectrogram using the mel filter bank approach [21]. We set the window size to 400 samples and use 64 mel filter banks. Finally, we compute the logarithm of the resulting spectrogram to obtain decibel (dB) values and normalize the spectrograms to the range $[-1,1]$ using the minimum and maximum log-spectrogram values over the entire training dataset.
III-C M2R2 Feature Extraction
Following common practices, we adopt a pretraining strategy to learn the M2R2 feature extractor, which is later used to extract multimodal features for training a TAS model [16]. Instead of using the full demonstration, we sample a shorter window of data from the recording. However, to ensure the model can still learn local relationships between actions, each window always contains data from three consecutive actions. For each action segment within the recording, we first obtain its start and end indices, denoted as $[i_{b},i_{e})$. To incorporate contextual information from neighboring actions, we expand these boundaries by a padding value $p$, which we experimentally set to $30$, which is enough time to observe meaningful states from neighboring action. This results in the extended boundaries $[i_{b}-p,i_{e}+p)$. To improve training efficiency, we randomly sample $N_{w}=100$ data points within these extended boundaries. The sampling is performed such that the number of selected samples within each section—$[i_{b}-p,i_{b})$, $[i_{b},i_{e})$, and $[i_{e},i_{e}+p)$—is proportional to the duration of that section. This ensures that each segment is adequately represented in the sampled window. Additionally, for each window, we return the natural language action descriptions of the actions present in the window in their respective order. We also provide a boundary sequence $B_{gt}\in\{0,1\}^{N_{w}}$, which marks where action transitions occur within the window. Since action transitions are typically smooth, we apply a Gaussian kernel to $B$ to obtain a soft boundary representation, $\bar{B}_{gt}=\lambda(B_{gt})$.
Next, our goal is to obtain the multimodal M2R2 feature $x_{i}$ for each sample within the window. To this end, we first compute a single representation for each sensor modality corresponding to sample $x_{i}$, which we then fuse using a self-attention modality transformer model [12]. For simplicity, we describe the procedure for a single sample, but this process is repeated for all sampled data points.
Depending on the modality, we adopt one of the following three strategies.
Figure 3: Temporal Fusion and Pretraining. Given a window $[i_{b}-p,i_{e}+p)$, we sample $N_{w}$ frames and extract features using our M2R2 feature extractor. A Temporal Fusion Transformer refines these features into $\widehat{X}$, which we average to obtain the window representation $E_{w}$. To learn action order, we minimize the distance between $E_{w}$ and a textual embedding $E_{s}$ generated from action labels by using a template. To enhance boundary detection, we minimize the MSE between the smoothed ground truth boundaries $\hat{B}_{gt}$ (obtained from the frame-wise dataset annotations) and predictions $B_{pred}$, obtained via a Boundary Regression Network.
Images: We use a Visual Transformer initialized with ActionCLIP weights [15], which produces the image representation $I_{i}\in\mathbb{R}^{D_{e}}$, where $D_{e}$ is the resulting M2R2 embedding size, which in our case is 512.
Audio: We use the Audio Spectrogram Transformer (AST) [21] to extract audio features from the log-mel spectrogram. As the audio representation, we take the classification token feature from the last hidden layer of the AST encoder, which produces the audio representation $A_{i}\in\mathbb{R}^{D_{e}}$.
Proprioceptive Sensor Readings: We employ a learnable temporal fusion approach to compute the representation of each proprioceptive sensor. After preprocessing, each proprioceptive sensor $s$ produces a time series of shape $D_{s}\times T_{s}$, where $D_{s}$ is the feature dimension of the sensor, and $T_{s}$ is the number of samples in the temporal dimension after resampling. To obtain a single representation for sensor $s$ at time step $i$, we first project the proprioceptive data into a higher-dimensional embedding space of size $D_{e}$ using a learned projection matrix $W_{p}^{s}$. We then apply a temporal fusion mechanism to capture relevant information along the time axis. Specifically, we use a learnable temporal embedding matrix $E_{t}^{s}\in\mathbb{R}^{D_{e}\times T_{s}}$, which is multiplied element-wise with the projected sensor features. Finally, we compute the mean over the temporal dimension to obtain a distinct representation for each sensor, denoted as $S_{i}^{s}\in\mathbb{R}^{D_{e}}$. Note that each sensor $s$ has its own unique projection and embedding matrices.
Finally, we fuse the information across all sensor modalities using a modality transformer followed by a multi-layer perceptron (MLP), akin to [24]. First, we construct a sequence of tokens by stacking the individual modality representations:
$\text{tok}=[I_{i},A_{i},S^{1}_{i},...,S^{s}_{i}]\in\mathbb{R}^{D_{e}\times(N_{s}+2)}.$ (2)
Next, we exchange information between modalities using a single self-attention transformer encoder layer [12]. This operation results in updated modality representations:
$\widehat{\text{tok}}=[\widehat{I_{i}},\widehat{A_{i}},\widehat{S^{1}_{i}},...,\widehat{S^{s}_{i}}].$ (3)
We then fuse all tokens by concatenating them into a single vector of shape $(1\times D_{e}\cdot(N_{s}+2))$, which is passed through an MLP to obtain the final multimodal representation: $x_{i}\in\mathbb{R}^{D_{e}}.$
III-D M2R2 Pretraining
We adopt a pretraining strategy similar to Br-Prompt [24]. To capture temporal dependencies between the multimodal features computed for each sampled datapoint in the window, we use a multi-head self-attention transformer encoder with $L$ layers, which we refer to as the Fusion Transformer. The Fusion Transformer takes as input the sequence of multimodal features $X\in\mathbb{R}^{T\times D_{e}}$ and produces a temporally enhanced sequence of features $\widehat{X}\in\mathbb{R}^{T\times D_{e}}$. To obtain a representation of the entire window, we compute the mean value over the temporal dimension, given by $E_{w}=\frac{1}{T}\sum_{i=1}^{T}\widehat{X}_{i}$, where $E_{w}\in\mathbb{R}^{D_{e}}$. Additionally, we input $\widehat{X}$ into an MLP, referred to as the Boundary Regression Network. The purpose of this network is to predict the probability of each sample within the window being a boundary sample. This encourages the model to learn more meaningful action boundaries from the underlying multimodal data. We denote the predicted boundary sequence as $B_{\text{pred}}\in\mathbb{R}^{T}$.
To pretrain the M2R2 feature extractor, we use two training objectives. First, we maximize the similarity of the window representation $E_{w}$ and an embedding $E_{s}$ of a sentence describing the order of actions (e.g., ”First, the robot does action A. Next, it performs action B. Finally, the machine executes action C”). We obtain this sentence embedding by first tokenizing the sentence and then using a transformer model initialized with ActionCLIP weights. Next, we maximize the similarity between the window embedding $E_{w}$ and its corresponding sentence embedding $E_{s}$, while minimizing the similarity to sentence embeddings of actions performed in different orders. To achieve this, we use the same loss function as presented in [16], which we denote as $\mathcal{L}_{\text{action}}$.
Additionally, to learn the correct boundaries, we minimize the average Mean Squared Error between the predicted and ground truth boundary sequences over the entire batch:
$\mathcal{L}_{\text{boundary}}=\frac{1}{B_{s}}\sum_{i=1}^{B_{s}}\frac{1}{T}\sum_{t=1}^{T}(\bar{B}_{gt}^{i}-B_{\text{pred}}^{i})^{2}.$ (4)
Our final loss is the sum of both loss functions:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{action}}+\mathcal{L}_{\text{boundary}}.$ (5)
TABLE I: Quantitative results. Best performing models in bold, second best are underlined.
Fine-grain Labels Coarse Label
Method F1@{10,25,50} Acc EDIT DR F1@{10,25,50} Acc EDIT DR
BR-Prompt [16]+DiffAct [11] 12.0 10.1 5.1 5.7 21.4 19.9 62.9 51.8 22.5 54.8 73.4 17.9
BOCPD [4] 39.2 32.6 12.8 35.0 19.7 17.4
AWE [1] 72.5 68.4 35.8 54.6 54.9 22.1
M2R2 (all) + MSTCN [8] 83,1 82,7 80,8 82,4 79,3 89,5 97,7 97,6 96,7 96,7 95,0 94,5
M2R2 (all) + ASRF [10] 83,5 83,4 82,4 82,7 82,5 95,1 98,1 98,1 97,1 96,0 96,9 95,9
M2R2 (all) + DiffAct [11] 78.1 77.7 74.6 74.9 68.7 81.4 95.1 95.0 92.1 91.7 90.3 83.9
IV Experiments
We conduct several experiments to evaluate the performance of the proposed M2R2 multimodal feature extractor. First, we compare the performance of the multimodal M2R2 features to several state-of-the-art temporal action segmentation approaches and video feature extractors. Next, we conduct extensive ablation studies to assess the impact of each modality and loss term on the robustness of the learned features.
IV-A Dataset
For our experiments, we use the REASSEMBLE [13] dataset, a contact-rich manipulation dataset where a robot performs assembly and disassembly tasks with various connectors, gears, pegs, and nuts. The dataset consists of 4,551 action demonstrations across 149 recordings. It presents significant challenges for temporal action segmentation due to the high visual similarity between objects (e.g., pegs and gears that differ only slightly in size) and the high median number of actions per recording (36). We select this dataset because it includes annotations for robot temporal action segmentation and provides measurements from multiple sensors. These include multi-view RGB cameras, multiple microphones, proprioceptive robot data, and force/torque readings at the robot’s wrist. For our experiments, we use both external cameras and the gripper-mounted microphone, along with force and torque measurements, the end-effector pose, and twist. The dataset contains 69 fine-grain, each combining an activity name (e.g., ”insert”) with an object name (e.g., ”USB”). We also define a coarse labeling scheme that uses only the activity names.
IV-B Metrics
The computer vision community evaluates temporal action segmentation (TAS) models using five standard metrics [11]: frame-level accuracy, EDIT score, and segmental F1 scores at 10%, 25%, and 50% overlap thresholds. Frame-level accuracy measures the percentage of correctly classified frames, while the EDIT score quantifies sequence alignment via the number of insertions, deletions, and swaps needed to match predictions to ground truth. Segmental F1 scores assess the quality of predicted action segments based on the minimum required overlap with ground truth.
Robotic TAS research employs similar metrics [4], often including the detection rate (DR) [4, 5], which evaluates action boundary detection. DR is computed as the F1 score by defining a ±$\pm t_{e}$-frame window around each ground truth boundary. A predicted boundary within this window counts as a true positive; multiple detections in the same window contribute one true positive, with the rest as false positives. Predictions outside any window are false positives, while unmatched ground truth boundaries are false negatives.
We report frame-level accuracy, EDIT score, F1 scores at 10%, 25%, and 50%, and DR for a comprehensive evaluation. We set $t_{e}=10$, corresponding to a 1 s time span.
IV-C Baselines
We evaluate the performance of BR-Prompt [16], a purely vision-based video feature extractor. We pretrain BR-Prompt using its default hyperparameters, sampling 16-frame windows during training. These windows typically span 2–3 actions and, at most, 6 actions. We compare our method to two state-of-the-art proprioception-based temporal action segmentation approaches used in robotics: BOCPD [4], which segments actions using the full 6DoF wrench, and AWE [1], which relies on position trajectories. We tune their hyperparameters to ensure the number of predicted segments aligns with the ground truth. Since these methods do not predict action labels, we assign each segment the ground truth class with the highest overlap. Additionally, we train three state-of-the-art deep learning models—MSTCN [8], ASRF [10], and DiffAct [11]—on M2R2 multimodal features.
IV-D Quantitative Evaluation
Table I presents the performance of each baseline model. We observe that deep learning-based TAS models trained with the multimodal M2R2 features achieve the best performance among all models. Specifically, ASRF achieves the highest performance with our features, attaining an F1@50 score of 82.4%, surpassing the unsupervised robotic TAS models, by at least 46.6 percentage points.
Among all baselines, the vision-only BR-Prompt video feature extractor performs the worst, with an F1@50 score of only 5.7%. We attribute this poor performance to BR-Prompt’s frame sampling strategy. In the REASSEMBLE dataset, long-duration actions like Insert and Remove are often separated by shorter actions such as Pick and Place. However, BR-Prompt’s window sampling strategy does not guarantee that frames from these short actions are included. If a short action is missed, two consecutive long actions may be grouped together, disrupting the expected action sequence in REASSEMBLE (e.g., Pick $\rightarrow$ Insert, Remove $\rightarrow$ Place). This can bias the model toward incorrect short-term dependencies between actions. In contrast, M2R2 ensures that no actions sampled in the window are omitted, and that each action is represented by a sufficient number of data points.
Figure 4: Quantitative evaluation of different baseline TAS models. AWE [1] performs poor in sections of highly nonlinear movement, while BOCPD [4] tends to over-segment in areas with high variation in force interactions (marked in black). The deep learning-based approaches occasionally misidentify objects.
Figure 5: Coarse level prediction compared to fine-grain level prediction. Same recording as in Figure 4.
The performance of deep learning-based TAS models trained with M2R2 features surpasses existing state-of-the-art unsupervised TAS approaches used in robotics. We only evaluate performance using the fine-grained label set, as the granularity of the labels does not impact the evaluation of the unsupervised approaches (where labels are assigned based on the overlap ground truth). During our experiments, we observed that unsupervised approaches tend to over-segment the demonstrations (e.g BOCPD and AWE), as shown in Figure 4.
IV-E Qualitative Evaluation
Figure 4 presents the qualitative evaluation results for some of the baseline models. As shown, the DiffAct model trained with BR-Prompt features struggles to segment both objects and actions in the recording. On the other hand, the BOCPD baseline over-segments the demonstration in areas where there is high variation in forces within the dataset. AWE over-segments in regions where highly nonlinear movements occur. The baseline models trained with M2R2 multimodal features tend to make mistakes and over-segment at certain points, but this is mostly due to object confusion, highlighting the challenging nature of the REASSEMBLE dataset. Figure 5 shows the coarse predictions of the same model on the same recording. As seen, the model correctly captures the ”remove” and ”place” actions.
TABLE II: Modality ablation study. All models where trained with the DiffAct TAS model [11]. We evaluate on fine-grain labels.
Method F1@{10,25,50} Acc EDIT DR
only vision 27,2 26,1 21,6 21,4 28,9 44.3
only audio 39,2 38,7 36,4 34,5 37,5 80.1
only prio 78,0 77,5 74,5 75,4 69,2 80.3
no gripper 67.9 67.2 64.2 65.1 59.2 82.2
no pos/vel 69.9 69.1 65.1 64.6 61.8 80.5
no F/T 72.9 72.4 69.9 70.4 65.2 80.6
all 78,1 77,7 74,6 74,9 68,7 82.4
IV-F Ablation Studies
We evaluate the impact of different sensory modalities on TAS performance by training the M2R2 feature extractor using only vision, only audio, or only proprioceptive data. Next, we combine both exteroceptive modalities and systematically remove specific sensors: pose and twist information, gripper width, or force and torque measurements.
Table II presents the performance of these sensory configurations. When using all modalities, we achieved the best performance. The worst performance occurs when using only vision, aligning with our previous observations where the BR-Prompt baseline also struggled. It seems due to the small sizes of objects which are easily occluded by the robot gripper. Audio, however, significantly improves the detection of action boundaries, achieving an 80.1% detection rate. This matches our expectations from analyzing the REASSEMBLE dataset, where action transitions are often marked by distinct sounds–for example, gripper closing when picking an object, opening when inserting, or characteristic sounds from connectors like an Ethernet cable or waterproof connector. Similarly, placing an object often results in an audible impact. However, segmentation performance remains poor, as identifying the manipulated object solely from sound is challenging, though some objects can still be recognized (see Figure 6).
Interestingly, TAS performance using only proprioceptive features is comparable to the model trained on all modalities. This may be due to the specificity of the REASSEMBLE dataset, where actions and objects can often be identified through proprioceptive cues alone. Pose and twist trajectories, gripper state, and force/torque measurements help distinguish both objects and activity types. However, due to the lack of diverse multimodal robotic datasets for TAS, we cannot further investigate this limitation. Future work should explore multimodal robotic TAS in more daily tasks with a higher object variability, such as meal preparation, where vision plays a crucial role in distinguishing similar objects (e.g., a bottle of orange juice vs. a bottle of water). One advantage of incorporating all sensory data is that it improves boundary placement within segmented demonstrations, as evidenced by the higher detection rate compare to a prioproceptive only model.
Finally, we analyze the impact of each proprioceptive measurement on TAS performance. We observe that removing gripper information has the most significant effect on the results. This is intuitive, as objects in the REASSEMBLE dataset can often be distinguished based on their size. Figure 6 illustrates cases where the absence of gripper information makes object identification more challenging. Similarly, removing pose and twist information also significantly affects model performance, as it makes action recognition more difficult. Additionally, removing force-torque (FT) data reduces action segmentation performance for similar reasons—it becomes harder to differentiate objects.
(a) Some object types can be determined from audio only.
(b) After removing gripper information some object categories and actions become hard to distinguish.
Figure 6: Exmaple predictions for different modality combinations used to train the M2R2 features.
V Conclusion
In this work, we propose M2R2, a multimodal robotic representation and pretraining strategy for temporal action segmentation (TAS). We leverage proprioceptive sensory information, including end-effector pose and twist, gripper width, and force-torque measurements, along with exteroceptive inputs such as vision and audio, to learn multimodal features for TAS. Unlike existing robotic multimodal TAS approaches, M2R2 features are model-agnostic, ensuring compatibility with various TAS architectures. We compare several state-of-the-art TAS models utilizing M2R2 features against leading visual and proprioception-based TAS approaches. All baselines trained with M2R2 features outperform prior robotics and computer vision methods by at least 46.6%. Additionally, we conduct an extensive ablation study to assess the impact of different modalities and sensors on TAS performance, showing that combining vision, audio, and proprioceptive data yields the highest performance.